# Word and Phrase Features in Graph Convolutional Network for Automatic Question Classification

### v2

@@ VI Conclusion
We develop and evaluate PQ-GCN, adapting GCNs to question classification by incorporating phrase features into the text graphs that can provide deeper semantic connections within questions. The proposed method shows promising classification performance compared to baseline models, capturing the intricate word and phrase relationships within question text. Question classification is an important task in building educational analytics systems, as we move towards an era of personalized learning, augmented with artificial intelligence. Finding suitable feature extraction methods is crucial in tackling this task, and we hope to provide sufficient basis for further research into innovative feature extraction methods for automatic text classification with graph neural networks.

### v3

@@ 8 Conclusion
We develop and evaluate PQ-GCN, which augments a SHINE-based GCN for question classification by incorporating phrase features into the text graphs that can provide deeper semantic connections within questions. The proposed method shows competitive classification performance compared to baseline models, especially in low-resource settings without the need for extensive pre-training. Question classification is an important task in building educational analytics systems, as we move towards an era of personalized learning with artificial intelligence. Finding suitable feature extraction methods is crucial in tackling downstream tasks in an increasingly data-scarce landscape, and our proposed method provides a parameter- and data-efficient approach to enhancing question classification.